Determine the de Broglie wavelength of a photon m
Answer #1

1. λ=h/Vm(proton)

h=6.63x10^-34J*s
mass(proton)=1.67265x10^-27 kg
v1=6.5x10^7

λ1=h/[V1]m(proton)=6.0981 e-15m

λ=h/Vm(electron)

h=6.63x10^-34J*s
mass(electron)=9.10938x10^-31 kg
v1=5.5x10^7

λ1=h/[V1]m(proton)=1.3233 e-15m

V = h/mλ = 3.4658 * 107